# Wireless Communication for Low-Altitude Economy with UAV Swarm Enabled Two-Level Movable Antenna System

### v3

@@ I Introduction
Low-altitude economy has emerged as a new integrated economic form, involving assorted low-altitude activities within sub-1,000-meter airspace domains [9, 3]. By leveraging manned/unmanned aerial vehicles (UAVs), and electric vertical take-off and landing (eVTOL) aircrafts, low-altitude economic ecosystems begin to flourish, which encompass a wide range of applications like logistics and delivery, agricultural plant protection, cultural tourism, environmental monitoring, and so on. In particular, benefiting from the advantages such as three-dimensional (3D) maneuverability, flexible deployment, and line-of-sight (LoS) air-to-air/ground communication links, UAV is regarded as an indispensable part of the low-altitude economy, which has driven various applications in wireless communication and sensing [47, 27, 37]. Specifically, from the perspective of wireless communication, there exist two paradigms, i.e., UAV-assisted wireless communication and cellular-connected UAV communication [46]. For example, UAV equipped with base station (BS) or relay can be deployed on demand to assist in the conventional cellular wireless networks, so as to satisfy the ubiquitous connectivity required by the future sixth-generation (6G) wireless network. Extensive research efforts have been devoted to this direction, including UAV channel and energy consumption modeling, performance analysis, and placement/trajectory design [47, 27]. On the other hand, the role of UAV can be shifted from the aerial BS/relay to aerial user equipments (UEs), thus enabling a new paradigm where both the ground and aerial UEs coexist in future wireless networks [46, 49, 47]. Moreover, as integrated sensing and communication (ISAC) has been identified as one of the six major usage scenarios for 6G [6], research on ISAC with UAV has gained a surge of interest recently [28, 30, 10]. Similarly, UAV can act as an aerial anchor for providing the ISAC service from the sky, or an aerial target to be sensed [37].
@@
However, UAVs are subjected to various practical constraints in terms of size, weight, dynamics, and power, rendering a single UAV challenging to execute sophisticated communication or sensing missions [47, 8, 11, 13, 48]. To tackle this issue, UAV swarm, i.e., a group of coordinated UAVs, can cooperatively accomplish sophisticated tasks, e.g., large-scale data collection and environment monitoring. Compared to single UAV, UAV swarm is able to significantly improve the overall payload capability and expand the communication and sensing coverage area. Moreover, the cooperative mechanism helps to reduce the task completion time and offers an enhanced robustness against anomalies or failures than single UAV, since the task of any malfunctioning UAV can be transferred to its neighboring functioning UAVs [7]. Given the above promising advantages, extensive research endeavours have been devoted to UAV swarm assisted communication and sensing. For example, in [5], an efficient channel estimation and self-positioning approach was proposed for UAV swarm communication with unknown displacements among UAVs. The authors in [26] proposed a 3D irregular terrain area coverage scheme with UAV swarm, and the results show that the proposed scheme can cover the area with little redundancy. Besides UAV swarm assisted communication, recent research works have studied the integration of ISAC with UAV swarm, since the LoS-dominating channel is particularly favorable for sensing [37]. Research on UAV swarm ISAC involves various performance metrics, including signal-to-interference-plus-noise ratio (SINR) and achievable rate for communication [3, 10], radar mutual information [17] and Cramér-Rao lower bound (CRLB) for sensing [30, 10, 40]. Despite the promising advantages, UAV swarm also involves several issues, such as the latency caused by information exchange among UAVs, complex collision avoidance planning, and high-precision localization and sensing for UAV swarm regulation [45, 8]. In particular, to tackle the challenge of accurately localizing and sensing the densely located swarm UAVs, a super-resolution communication and sensing method was proposed for UAV swarm in [45].
@@ II System Model And Problem Formulation
where ${\alpha_{k}}=\sqrt{{\beta_{0}}}/{r_{k}}$ denotes the channel gain of UE $k$, with $\beta_{0}$} being the channel power at the reference distance of ${d_{0}}=1$ m. Besides, ${{\bf{a}}_{k}}\left({\left\{{{{\bf{\bar{q}}}_{l,m}},\forall l,m}\right\}}\right)\in{\mathbb{C}}^{LM\times 1}$ denotes the receive array response vector of UE $k$, given by
@@ III-A Single UE and Two UEs
Next, we consider the classic MRC, zero-forcing (ZF) and minimum mean-square error (MMSE) beamforming schemes for the case of two UEs. Denote by ${\xi_{k,k^{\prime}}}\triangleq\frac{{{{\left|{{\bf{h}}_{k}^{H}{{\bf{h}}_{k^{\prime}}}}\right|}^{2}}}}{{{{\left\|{{{\bf{h}}_{k}}}\right\|}^{2}}{{\left\|{{{\bf{h}}_{k^{\prime}}}}\right\|}^{2}}}}$ the channel’s squared-correlation coefficient between UEs $k$ and $k^{\prime}$, where $k,k^{\prime}=1,2$ and $k\neq k^{\prime}$. Under the far-field UPW model, we have ${\left\|{{{\bf{h}}_{k}}}\right\|^{2}}={\left|{{\alpha_{k}}}\right|^{2}}L$. The resulting SINR/SNR of UE $k$ after applying the three beamforming schemes are respectively given by [23, 18]
@@
To tackle problem (19), we first study the property of the objective function. Note that when the movable region size is much smaller than the link distance, the UPW model gives a valid approximation of spherical wavefront for channels between UAVs and ground UEs [18], which may correspond to the cases of small movable region and/or high UAV operation altitude. In this case, the receive response vector in (7) reduces to ${{\bf{a}}_{k}}={e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{r_{k}}}}{[{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{k}}\left({{{\bf{q}}_{1}}}\right)}},\cdots,{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\varphi_{k}}\left({{{\bf{q}}_{L}}}\right)}}}]^{T}}$, where ${\varphi_{k}}\left({{{\bf{q}}_{l}}}\right)={\bm{\kappa}}_{k}^{T}{{\bf{q}}_{l}}$. By substituting it into the objective function, we have
$\frac{1}{{{L^{2}}}}{\left|{{\bf{a}}_{k}^{H}{{\bf{a}}_{k^{\prime}}}}\right|^{2}}$ $=\frac{1}{{{L^{2}}}}{\left|{\sum\limits_{l=1}^{L}{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}\left({{\varphi_{k}}\left({{{\bf{q}}_{l}}}\right)-{\varphi_{k^{\prime}}}\left({{{\bf{q}}_{l}}}\right)}\right)}}}}\right|^{2}}$ (20)
$=\frac{1}{{{L^{2}}}}{\left|{\sum\limits_{l=1}^{L}{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}{{\left({{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right)}^{T}}{{\bf{q}}_{l}}}}}}\right|^{2}}.$
@@
where ${\bf{q}}_{1}$} can be any vector to guarantee ${{\bf{q}}_{l}}\in{\cal C}$, $\forall l$, and $\varsigma^{\star}$} is given by
@@
Note that Theorem 1 extends the 1D result in [50] to the general 3D placement, and the derived closed-form solution $\varsigma^{\star}$} can be a fraction, which removes the requirement of being an integer imposed in [50]. Moreover, UAV placement position ${{\bf{q}}_{l}}\in{\cal C}$, $\forall l$, can be satisfied when $\left\|{{{\bf{q}}_{L}}-{{\bf{q}}_{1}}}\right\|=\frac{{\left({L-1}\right)\left({{\varsigma^{\star}}+1/L}\right)\lambda}}{{\left\|{{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right\|}}$ is no greater than the maximum distance between any two points in $\cal C$}.
With (21) and (22), we have ${\xi_{k,k^{\prime}}}=0$, and ${\gamma_{{\rm{MRC}},k}}={\gamma_{{\rm{ZF}},k}}\ ={\gamma_{{\rm{MMSE}},k}}={{\bar{P}}_{k}}{\left\|{{{\bf{h}}_{k}}}\right\|^{2}}={{\bar{P}}_{k}}{\left|{{\alpha_{k}}}\right|^{2}}L$, i.e., they are identical to the single UE SNR without IUI. Thus, an IUI-free communication can be achieved with the UAV swarm enabled MA system, by setting the UAVs’ static positions according to Theorem 1. It is worth noting that the inter-UAV distance is $\frac{{\left|{{\varsigma^{\star}}+1/L}\right|\lambda}}{{\left\|{{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right\|}}\geq d_{\min}$, where the minimum distance $d_{\min}$ is in general much larger than half wavelength. Thus, the UAV swarm forms a USA for the IUI-free communication. In particular, compared to the classic compact array where antenna spacing is separated by half wavelength, USA can achieve a larger array aperture, and thus a narrower beamwidth of the main lobe is enabled.
@@ III-B1 Optimization of $\bf V$} With Given $\bf Q$}
where ${\delta_{k,i,l}}=\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|}\sqrt{2+{{\left({\frac{{2\pi}}{\lambda}}\right)}^{2}}{{\left({1+\frac{{{C_{\max}}}}{{r_{i}^{2}}}}\right)}^{2}}}$, we have ${\nabla^{2}}{g_{k,i,l}}\preceq{\delta_{k,i,l}}{\bf{I}}$, ${\bf{q}}_{l}^{\left(j\right)}$} denotes the resulting placement position of UAV $l$ in the $j$-th iteration, and $\nabla{g_{k,i,l}}({{\bf{q}}_{l}^{\left(j\right)}})$ denotes the gradient of $g_{k,i,l}$}} over ${\bf{q}}_{l}^{\left(j\right)}$.
@@ IV-B1 Optimization of $\bf V$} With Given $\bf Q$} and $\bf B$}
where ${{\bf{d}}_{i,l}}={[{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{k,l}^{T}{{\bf{b}}_{l,1}}}},\cdots,{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{k,l}^{T}{{\bf{b}}_{l,M}}}}]^{T}}$, ${{\bf{V}}_{k,l,l^{\prime}}}\in{{\mathbb{C}}^{M\times M}}\triangleq{{\bf{v}}_{k,l}}{\bf{v}}_{k,l^{\prime}}^{H}$, and ${{{\bf{\tilde{V}}}}_{k,l,l}}\in{{\mathbb{C}}^{\left({L-1}\right)M\times\left({L-1}\right)M}}\triangleq{{{\bf{\tilde{v}}}}_{k,l}}{\bf{\tilde{v}}}_{k,l}^{H}$. Moreover, ${\hat{g}}_{k,i,l}$}} is a constant term, and ${\tilde{g}}_{k,i,l}$}} can be expressed in terms of ${\bf{q}}_{l}$}} as
@@
$\mathop{\max}\limits_{{b_{l,m}}}$ $\ \mathop{\min}\limits_{k\in{\cal K}}\ {{R_{k}}}$ (49)
${\rm{s.t.}}$ $\ {b_{l,m}}\in{\tilde{\cal C}}$,
@@
Let ${{\bf{\tilde{h}}}_{i,l,m}}\in{{\mathbb{C}}^{\left({LM-1}\right)\times 1}}$ and ${{\bf{\tilde{v}}}_{k,l,m}}\in{{\mathbb{C}}^{\left({LM-1}\right)\times 1}}$ denote the resulting channel and beamforming vector after removing $h_{i,l,m}$} and $v_{i,l,m}$}, respectively, where $h_{i,l,m}$} and $v_{i,l,m}$} are the $\left({\left({l-1}\right)M+m}\right)$-th elements of ${\bf{h}}_{i}$}} and ${\bf{v}}_{k}$}}, respectively. The term ${f_{k,i}}={\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}}\right|^{2}}$ can be further expressed as
@@
The main procedures for solving problem (P1) are summarized in Algorithm 2. The complexity for obtaining the receive beamforming in step 3 is ${\cal O}({K{(LM)^{3}}})$. From step 4 to step 6, the complexity is approximately ${\cal O}({I_{1}}L{{({2K})}^{3}})$, where $I_{1}$ denotes the maximum number of iterations required by SCA for convergence. The complexity from step 7 to step 11 is approximately given by ${\cal O}({I_{2}}L(M-1){{({2K})}^{3}})$, where $I_{2}$ denotes the maximum number of iterations to converge required by step 9. As a result, the total computational complexity of Algorithm 2 is ${\cal O}({{I_{3}}K{(LM)^{3}}+{I_{2}}{I_{3}}L(M-1){({2K})}^{3}})$, where $I_{3}$ denotes the number of iterations required by the alternating optimization for convergence.

### v4

@@ I Introduction
Low-altitude economy has emerged as a new integrated economic form, involving assorted low-altitude activities within sub-1,000-meter airspace domains [9, 3]. By leveraging manned/unmanned aerial vehicles (UAVs), and electric vertical take-off and landing (eVTOL) aircrafts, low-altitude economic ecosystems begin to flourish, which encompass a wide range of applications like logistics and delivery, agricultural plant protection, cultural tourism, environmental monitoring, and so on. In particular, benefiting from the advantages such as three-dimensional (3D) maneuverability, flexible deployment, and line-of-sight (LoS) air-to-air/ground communication links, UAV is regarded as an indispensable part of the low-altitude economy, which has driven various applications in wireless communication and sensing [47, 27, 37]. Specifically, from the perspective of wireless communication, there exist two paradigms, i.e., UAV-assisted wireless communication and cellular-connected UAV communication [46]. For example, UAV equipped with a base station (BS) or relay can be deployed on demand to assist in the conventional cellular wireless networks, so as to satisfy the ubiquitous connectivity required by the future sixth-generation (6G) wireless network. Extensive research efforts have been devoted to this direction, including UAV channel and energy consumption modeling, performance analysis, and placement/trajectory design [47, 27]. On the other hand, the role of UAV can be shifted from the aerial BS/relay to aerial user equipments (UEs), thus enabling a new paradigm where both the ground and aerial UEs coexist in future wireless networks [46, 49, 47]. Moreover, as integrated sensing and communication (ISAC) has been identified as one of the six major usage scenarios for 6G [6], research on ISAC with UAV has gained a surge of interest recently [28, 30, 10]. Similarly, UAV can act as an aerial anchor for providing the ISAC service from the sky, or an aerial target to be sensed [37].
@@
However, UAVs are subject to various practical constraints in terms of size, weight, dynamics, and power, rendering a single UAV challenging to execute sophisticated communication or sensing missions [47, 8, 11, 13, 48]. To tackle this issue, UAV swarm, i.e., a group of coordinated UAVs, can cooperatively accomplish sophisticated tasks, e.g., large-scale data collection and environment monitoring. Compared to single UAV, UAV swarm is able to significantly improve the overall payload capability and expand the communication and sensing coverage area. Moreover, the cooperative mechanism helps to reduce the task completion time and offers an enhanced robustness against anomalies or failures than single UAV, since the task of any malfunctioning UAV can be transferred to its neighboring functioning UAVs [7]. Given the above promising advantages, extensive research endeavours have been devoted to UAV swarm assisted communication and sensing. For example, in [5], an efficient channel estimation and self-positioning approach was proposed for UAV swarm communication with unknown displacements among UAVs. The authors in [26] proposed a 3D irregular terrain area coverage scheme with UAV swarm, and the results show that the proposed scheme can cover the area with little redundancy. Besides UAV swarm assisted communication, recent research works have studied the integration of ISAC with UAV swarm, since the LoS-dominating channel is particularly favorable for sensing [37]. Research on UAV swarm ISAC involves various performance metrics, including signal-to-interference-plus-noise ratio (SINR) and achievable rate for communication [3, 10], radar mutual information [17] and Cramér-Rao lower bound (CRLB) for sensing [30, 10, 40]. Despite the promising advantages, UAV swarm also involves several issues, such as the latency caused by information exchange among UAVs, complex collision avoidance planning, and high-precision localization and sensing for UAV swarm regulation [45, 8]. In particular, to tackle the challenge of accurately localizing and sensing the densely located swarm UAVs, a super-resolution communication and sensing method was proposed for UAV swarm in [45].
@@ II System Model And Problem Formulation
where ${\alpha_{k}}=\sqrt{{\beta_{0}}}/{r_{k}}$, with $\beta_{0}$} being the channel power at the reference distance of ${d_{0}}=1$ m. Besides, ${{\bf{a}}_{k}}\left({\left\{{{{\bf{\bar{q}}}_{l,m}},\forall l,m}\right\}}\right)\in{\mathbb{C}}^{LM\times 1}$ denotes the receive array response vector of UE $k$, given by
@@ III-A Single UE and Two UEs
Next, we consider the classic MRC, zero-forcing (ZF) and minimum mean-square error (MMSE) beamforming schemes for the case of two UEs. Denote by ${\xi_{k,k^{\prime}}}\triangleq\frac{{{{\left|{{\bf{h}}_{k}^{H}{{\bf{h}}_{k^{\prime}}}}\right|}^{2}}}}{{{{\left\|{{{\bf{h}}_{k}}}\right\|}^{2}}{{\left\|{{{\bf{h}}_{k^{\prime}}}}\right\|}^{2}}}}$ the channel’s squared-correlation coefficient between UEs $k$ and $k^{\prime}$, where $k,k^{\prime}=1,2$ and $k\neq k^{\prime}$. Besides, we have ${\left\|{{{\bf{h}}_{k}}}\right\|^{2}}={\left|{{\alpha_{k}}}\right|^{2}}L$. The resulting SINR/SNR of UE $k$ after applying the three beamforming schemes are respectively given by [23, 18]
@@
To tackle problem (19), we first study the property of the objective function. Note that when the movable region size is much smaller than the link distance, the UPW model gives a valid approximation of spherical wavefront for channels between UAVs and ground UEs [18], which may correspond to the cases of small movable region and/or high UAV operation altitude. In this case, the receive response vector in (7) reduces to ${{\bf{a}}_{k}}={e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{r_{k}}}}{[{{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{k}^{T}{{\bf{q}}_{1}}}},\cdots,{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{k}^{T}{{\bf{q}}_{L}}}}}]^{T}}$. By substituting it into the objective function, we have
$\frac{1}{{{L^{2}}}}{\left|{{\bf{a}}_{k}^{H}{{\bf{a}}_{k^{\prime}}}}\right|^{2}}=\frac{1}{{{L^{2}}}}{\left|{\sum\limits_{l=1}^{L}{{e^{{\rm{j}}\frac{{2\pi}}{\lambda}{{\left({{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right)}^{T}}{{\bf{q}}_{l}}}}}}\right|^{2}}.$ (20)
@@
where ${\bf{q}}_{1}$} can be any vector to guarantee ${{\bf{q}}_{r}}+{{\bf{q}}_{l}}\in{\cal C}$, $\forall l$, and $\varsigma^{\star}$} is given by
@@
Note that Theorem 1 extends the 1D result in [50] to the general 3D placement, and the derived closed-form solution $\varsigma^{\star}$} can be a fraction, which removes the requirement of being an integer imposed in [50]. Moreover, a sufficient condition for ${{\bf{q}}_{r}}+{{\bf{q}}_{l}}\in{\cal C}$, $\forall l$, is that $\cal C$} contains a line segment parallel to $({\bm{\kappa}}_{k}-{\bm{\kappa}}_{k^{\prime}})$ with length no smaller than $\left\|{{{\bf{q}}_{L}}-{{\bf{q}}_{1}}}\right\|=\frac{{\left({L-1}\right)\left({{\varsigma^{\star}}+1/L}\right)\lambda}}{{\left\|{{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right\|}}$.
With (21) and (22), we have ${\xi_{k,k^{\prime}}}=0$, and ${\gamma_{{\rm{MRC}},k}}={\gamma_{{\rm{ZF}},k}}\ ={\gamma_{{\rm{MMSE}},k}}={{\bar{P}}_{k}}{\left\|{{{\bf{h}}_{k}}}\right\|^{2}}={{\bar{P}}_{k}}{\left|{{\alpha_{k}}}\right|^{2}}L$, i.e., they are identical to the single UE SNR without IUI. Thus, an IUI-free communication can be achieved with the UAV swarm enabled MA system, by setting the UAVs’ static positions according to Theorem 1. It is worth noting that the adjacent inter-UAV distance is $\frac{{\left|{{\varsigma^{\star}}+1/L}\right|\lambda}}{{\left\|{{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right\|}}\geq d_{\min}$, where the minimum distance $d_{\min}$ is in general much larger than half wavelength. Thus, the UAV swarm forms a USA for the IUI-free communication. In particular, compared to the classic compact array where antenna spacing is separated by half wavelength, USA can achieve a larger array aperture, and thus a narrower beamwidth of the main lobe is enabled.
@@ III-B1 Optimization of $\bf V$} With Given $\bf Q$}
where ${\delta_{k,i,l}}=\frac{{2\pi}}{\lambda}\sum\limits_{l^{\prime}=1,l^{\prime}\neq l}^{L}{2{{\left|{{\alpha_{i}}}\right|}^{2}}\left|{{V_{k,l,l^{\prime}}}}\right|}\sqrt{\frac{2}{{r_{i}^{2}}}+{{\left({\frac{{2\pi}}{\lambda}}\right)}^{2}}{{\left({1+\frac{{{C_{\max}}}}{{r_{i}^{2}}}}\right)}^{2}}}$, ${\bf{q}}_{l}^{\left(j\right)}$} denotes the resulting placement position of UAV $l$ in the $j$-th iteration, and $\nabla{g_{k,i,l}}({{\bf{q}}_{l}^{\left(j\right)}})$ denotes the gradient of $g_{k,i,l}$}} over ${\bf{q}}_{l}^{\left(j\right)}$.
@@ IV-B1 Optimization of $\bf V$} With Given $\bf Q$} and $\bf B$}
where ${{\bf{d}}_{i,l}}={[{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{i,l}^{T}{{\bf{b}}_{l,1}}}},\cdots,{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{i,l}^{T}{{\bf{b}}_{l,M}}}}]^{T}}$, ${{\bf{V}}_{k,l,l^{\prime}}}\in{{\mathbb{C}}^{M\times M}}\triangleq{{\bf{v}}_{k,l}}{\bf{v}}_{k,l^{\prime}}^{H}$, and ${{{\bf{\tilde{V}}}}_{k,l,l}}\in{{\mathbb{C}}^{\left({L-1}\right)M\times\left({L-1}\right)M}}\triangleq{{{\bf{\tilde{v}}}}_{k,l}}{\bf{\tilde{v}}}_{k,l}^{H}$. Moreover, we adopt the quasi-static local AoA approximation, ${\hat{g}}_{k,i,l}$}} is a constant term, and ${\tilde{g}}_{k,i,l}$}} can be expressed in terms of ${\bf{q}}_{l}$}} as
@@
$\mathop{\max}\limits_{{b_{l,m}},{\rho}}$ $\ \rho$ (49)
${\rm{s.t.}}$ $\ {R_{k}}\geq\rho,\ \forall k,$
$\ {b_{l,m}}\in{\tilde{\cal C}}$,
@@
Let ${{\bf{\tilde{h}}}_{i,l,m}}\in{{\mathbb{C}}^{\left({LM-1}\right)\times 1}}$ and ${{\bf{\tilde{v}}}_{k,l,m}}\in{{\mathbb{C}}^{\left({LM-1}\right)\times 1}}$ denote the resulting channel and beamforming vector after removing $h_{i,l,m}$} and $v_{i,l,m}$}, respectively, where $h_{i,l,m}$} and $v_{k,l,m}$} are the $\left({\left({l-1}\right)M+m}\right)$-th elements of ${\bf{h}}_{i}$}} and ${\bf{v}}_{k}$}}, respectively. The term ${f_{k,i}}={\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}}\right|^{2}}$ can be further expressed as
@@
The main procedures for solving problem (P1) are summarized in Algorithm 2. The complexity for obtaining the receive beamforming in step 3 is ${\cal O}({K{(LM)^{3}}})$. From step 4 to step 6, the complexity is approximately ${\cal O}({I_{1}}L{{({2K})}^{3}})$, where $I_{1}$ denotes the maximum number of iterations required by SCA for convergence. The complexity from step 7 to step 11 is approximately given by ${\cal O}({I_{2}}L(M-1){{({2K})}^{3}})$, where $I_{2}$ denotes the maximum number of iterations to converge required by step 9. As a result, the total computational complexity of Algorithm 2 is ${\cal O}({{I_{3}}K{(LM)^{3}}+{I_{1}}{I_{3}}L{{({2K})}^{3}}+{I_{2}}{I_{3}}L(M-1){({2K})}^{3}})$, where $I_{3}$ denotes the number of iterations required by the alternating optimization for convergence.